5. , which of the following statements is (are) correct?
(x) Transactions involving items produced in the past are included in current GDP because GDP measures the value of goods and services sold in the current year.
(y) Robert, a United States citizen, works only in Mexico. The value added to production from his employment is included in the GDP of the United States, but not Mexican GDP.
(z) GDP excludes items produced and sold illegally and items that are produced and consumed at home.
. A. (x), (y) and (z)

B. (x) and (y) only

C. (x) and (z) only

D. (y) and (z) only
E. (z) only

which of the following statements is (are) correct?
(x) If nominal GDP is $16 trillion and real GDP is $18 trillion, then the GDP deflator is 88.9.
(y) If the GDP deflator is 150 and nominal GDP is $480 billion, then real GDP is $320 billion.
(z) If a country has real GDP of $500 billion and GDP deflator of 80, then its nominal GDP is $625 billion
A. (x), (y) and (z) B. (x) and (y) only
C. (x) and (z) only D. (y) and (z) only
E. (x) only
Answer #1

5. (x) false only that production is included in current gdp which produced in current year only

(y) false The value added to production from his employment is included in the GDP of Mexica in the form of net factor income to abroad.

(z) true illegally items produced are not counted in calculation of gdp

E. (z) only

6. GDP Deflator = Nominal GDP/Real GDP

(x) true gdp deflator = 16/18 = 88.89%

(y) true real gdp = 480/150% = $320

(z) false nominal gdp = 500*80% =$400

B. x and y only